A random sample of n = 45 values give = 63. Can this sample be regarded as drawn from a


A normal population with mean μ = 70 and s2 =25?


1
Expert's answer
2021-01-13T19:31:49-0500

Hypothesis

"H_0:\bar x =70"

"H_1:\bar x \not=70"

Since sample size "45 >30" we use the normal distribution table



"Z= \frac{\bar x - \mu}{\sqrt{s^2}}"


"\bar x= 63"

"\mu=70"

"s^2= 25"


Therefore


"Z= \frac{63 - 70}{\sqrt{25}}= -1.4"

At a "95\%" level of confidence, the sample mean can be concluded to be equal to the population mean "\mu =70" if:



"-1.96<Z<1.96"

Since the Z value obtained is within the range i.e.


"-1.96<-1.4<1.96"

We conclude that the sample is from the normal population with mean = 70 and variance = 25